OpenStudy (anonymous):

Identify which distributive property problem below would help you solve the following scenario: Your Math class has 32 students. You would like to treat the class to cookies. You would like each person to have 5 cookies. Which expression would be a quick way to figure out how many cookies you would need to bake? a.) 5(32 + 2) b.) 5(30 + 2) c.) 5(32 − 2) d.) 5(30 − 2)

OpenStudy (cwrw238):

you'll need 5*32 cookies = 160 which one of the four will give you 160?

OpenStudy (cwrw238):

its not d 5(30 - 2) = 5*30 - 5*2 = 150 -10 = 140